OpenStudy (anonymous):

How would I do this problem? (13 + radical 5) + (2 + 3 radical 5)

OpenStudy (anonymous):

\[(13+\sqrt{5}) + (2+3\sqrt{5})\]

OpenStudy (anonymous):

imagine \[\sqrt{5}=x\] then add like terms

OpenStudy (anonymous):

13+x+2+3x

OpenStudy (anonymous):

so i would add x and 3x to get 4x which is actually \[4\sqrt{5}\]

OpenStudy (anonymous):

and then add 13 and 2? so my final answer would be\[14+4\sqrt{5}\] ?

OpenStudy (anonymous):

its \[15+4\sqrt{5}\]

OpenStudy (anonymous):

that's what i meant haha! typo! :-)

OpenStudy (anonymous):

we have touse calculator or else approaximate it

OpenStudy (anonymous):

yeah i meant 15 instead of 14, thank you!